ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 84, 1993

Ash, J. M., E. Rieders and R. P. Kaufman The Cantor–Lebesgue property

179

Ballester-Bolinches, A., K. Doerk and L. M. Ezquerro A question in the theory of saturated formations of finite soluble groups

229

Bonnet, R. and S. Shelah On HCO spaces. An uncountable compact T_{2} space, different from ℵ_{1} +1, which is homeomorphic to each of its uncountable closed subspaces

289

Bourgain, J. On the spectral type of Ornstein’s class one transformations

53

Bramson, M. and S. Kalikow Nonuniqueness in g-functions  153

Carothers, N. L., R. Haydon and P.-K. Lin On the isometries of the Lorentz function spaces

265

Defant, M. and M. Junge Some estimates on entropy numbers  417

Doerk, K. See A. Ballester-Bolinches

Ezquerro, L. M. See A. Ballester-Bolinches

Fuchs, L. and M. Magidor Butler groups of arbitrary cardinality  239

Haydon, R. See N. L. Carothers

Hitczenko, P. Domination inequality for martingale transforms of a Rademacher sequence

161

Jin, R. See S. Shelah

Judah, H. and S. Shelah Baire property and axiom of choice  435

Junge, M. See M. Defant

Kalikow, S. See M. Bramson

Kano, H. and I. Shiokawa Rings of normal and nonnormal numbers 403
Katok, S. and P. Sarnak Heegner points, cycles and Maass forms 193

Kaufman, R. P. See J. M. Ash

Lima, Å. The metric approximation property, norm-one projections and intersection properties of balls

451

Lin, P.-K. See N. L. Carothers

Ling, S. The old subvariety of J_{0}(pq) and the Eisenstein kernel in Jacobians 365 Lluch, A. and V. Miquel Comparing the relative volume with a

revolution manifold as a model

385

Lochter, M. On equivalent number fields with special Galois groups  89

Magidor, M. See L. Fuchs

Mastrangelo, L. B. On radicals of triangular operator algebras  33

Miquel, V. See A. Lluch Moriah, Y. and V. Shpilrain Non-tame automorphisms of exten-

sions of periodic groups

17

Odell, E., H. P. Rosenthal and Th. Schlumprecht On weakly null FDD’s in Banach spaces

333

Quinto, E. T. Pompeiu transforms on geodesic spheres in real analytic manifolds

353

Rieders, E. See J. M. Ash Rosenthal, H. P. See E. Odell Sarnak, P. See S. Katok Schlumprecht, Th. See E. Odell

Schramm, O. Square tilings with prescribed combinatorics  97

Schwarz, B. Carathéodory balls and norm balls of the domain H =

{(z_{1}, z_{2}) ∈ C^{2} : |z_{1}| + |z_{2}| ≤ 1} 119

Shelah, S. and R. Jin A model in which there are Jech–Kunen trees but there are no Kurepa trees

1

Shelah, S. See R. Bonnet Shelah, S. See H. Judah Shiokawa, I. See H. Kano Shpilrain, V. See Y. Moriah

Stöhr, R. Homology of free Lie powers and torsion in groups  65
Zeitz, P. The centralizer of a rank-one flow  129
Zisser, I. Irreducible products of characters in A_{n}  147